# Microbenchmarking NVIDIA’s Blackwell Architecture: An in-depth Architectural Analysis

### v1

@@ VII-B2 Batch Size Sensitivity
| STREAM Triad | BW (TB/s) | 7.48 | 4.38 | 1.71 | HBM3e |
| SpMV (compressed) | GFLOPS | 5.04 | 3.2 | 1.58 | Decomp engine |
| GPT Training (1.3B) | tok/s | 14,397 | 9,240 | 1.56 | CTA pairs, TMEM, TC |
| ResNet Training | img/s | 2,436 | 1,580 | 1.54 | 5th Gen TC, mem BW |
| Energy Eff. (Training) | tok/s/W | 22.2 | 15.6 | 1.42 | Process, efficiency |
@@ VII-C1 FP64 Performance
Scientific computing applications present fundamentally different computational characteristics compared to deep learning workloads, needing high-precision arithmetic, sustainable memory bandwidth, and irregular access patterns. We evaluate dense matrix multiplication (DGEMM) performance using dobule-precision FP arithmetic (FP64), which remains essential for scientific simulations requiring numerical accuracy. Table XII presents our results across varying matrix dimensions. The B200 achieves 36.3 TFLOPS at large matrix size, representing 80.7% utilization of its 40 TFLOPS theoretical peak [nvidia_blackwell_b200_datasheet], compared to the H200’s 18.9 TFLOPS (55.6% of 34 TFLOPS). The additional 45% efficiency improvement results from TMEM-enabled accumulation that reduces L2 cache traffic and improved memory access coalescing patterns that better utilize the available memory bandwidth.
@@ VII-C2 Sustained Memory Bandwidth
Memory-intensive scientific applications require sustained high-bandwidth data movement capabilities. We employ the STREAM Triad benchmark to measure achievable memory banwidth across different working set sizes, as shown in Table XIII.
| Array Size | B200 (TB/s) | H200 (TB/s) | B200 % | H200 % |
| 4GB | 4.134 | 2.88 | 51.7 | 60.0 |
| 16GB | 4.141 | 2.91 | 51.8 | 60.6 |
| 64GB | 7.42 | 4.35 | 92.8 | 90.6 |
| 128GB | 7.48 | 4.38 | 93.5 | 91.3 |
The results show, smaller arrays that git within the GPU’s cache hierarchy achieve lower absolute bandwidth but higher efficiency on the H200 (60-60.6%) compared to the B200 (51.7-51.8%). This reflects the H200’s is more tuned for smaller working sets. Though, for working sets that exceed cache capacity, both architectures achieve excellent memory bandwidth utilization exceeding 90%. The 1.71 speedup closely matches raw bandwidth ratio (8.0/4.8), demonstrating memory-bound codes scale linearly with available bandwidth.
@@ VII-C3 Sparse Operations
| webbase-1M | 99.99% | 5.09 | 3.16× | 39.32 |
| circuit5M | 99.95% | 4.96 | 3.16× | 201.44 |
| ldoor | 99.98% | 5.04 | 3.16× | 71.93 |
The decompression engine yields consistent 3.16 speedups on sparse matrices. Run-length encoding (RLE) achieves 8.2 compression ratio for sparse row pointer arrays. The dedicated decompression hardware introduces less than 5% latency overhead while providing 35% traffic reduction in memory traffic for pointer-intensive workloads.
@@ VII-D Mixed-Precision Training: End-to-End Training Performance
We present comprehensive training benchmarks across different model architectures to assess the practical impact of architectural improvements in realistic training scenarios, as summarized in Table XI. Training achieves consistent 1.54–1.56 speedup, decomposing into SM count (1.09), CTA pairing (1.27), and TMEM (1.26). Energy efficiency improves 42% for GPT training despite 14% higher power consumption.
@@ VIII Discussion
Table X provides a comprehensive overview of performance improvements across all evaluated workload categories, highlighting the specific architectural features responsible for each performance gain. Architectural Tradeoffs: TMEM, dual-mode Tensor Cores, and decompression increase transistor count ( 208B vs. 180B) but deliver 1.5–3.9 gains. The 256KB TMEM per SM (10% of SM memory) achieves 61–82% hit rates, validating sizing decisions. Software Ecosystem: CUDA 13.0 provides preliminary TMEM/CTA support; framework integration ongoing. FP6 hardware support exists but lacks software tooling. FP4/FP6 require per-layer precision selection—8.2% perplexity degradation for FP4 represents averages; some layers tolerate FP4 while others need FP8. Deployment: For LLM inference, B200 provides 1.8–3.9 advantages; FP4 practical for 70B models. Training improvements (1.54–1.56) enable 33% larger batches. HPC gains (1.92 FP64) competitive for scientific computing.
@@ IX Conclusion
NVIDIA’s B200 GPU marks a major shift in GPU architectures. Our work presents the first detailed microbenchmark suite-based characterization of the NVIDIA Blackwell B200 GPU. Our work offers insights into its architectural innovations and performance behavior. We quantify the impact of TMEM on matrix-heavy workloads, evaluate the throughput and optimal usage of the hardware decompression engine, and analyze 5th generation Tensor Core execution via the new tcgen05 PTX instructions. Our study further assesses FP4 and FP6 precision trade-offs, benchmarks Blackwell across diverse workloads—including LLM inference, scientific kernels, and mixed-precision training—and distills actionable performance guidelines for developers targeting this next-generation architecture.

### v2

@@ VII-B2 Batch Size Sensitivity
| STREAM Triad (4–16 GB) | BW (TB/s) | 4.14 | – | – | HBM3e (B200 only) |
| SpMV (compressed) | GFLOPS | 5.08 | 3.2 | 1.58 | Decomp engine |
| GPT Training (1.3B) | tok/s | 14,363 | 9,240 | 1.55 | CTA pairs, TMEM, TC |
| ResNet Training | img/s | 2,928 | 1,580 | 1.85 | 5th Gen TC, mem BW |
| Energy Eff. (Training) | tok/s/W | 20.63 | 15.6 | 1.32 | Process, efficiency |
@@ VII-C1 FP64 Performance
Scientific computing applications present fundamentally different computational characteristics compared to deep learning workloads, needing high-precision arithmetic, sustainable memory bandwidth, and irregular access patterns. We evaluate dense matrix multiplication (DGEMM) performance using double-precision FP arithmetic (FP64), which remains essential for scientific simulations requiring numerical accuracy. Table XII presents our results across varying matrix dimensions. The B200 achieves 36.3 TFLOPS at large matrix size, representing 80.7% utilization of its 40 TFLOPS theoretical peak [2], compared to the H200’s 18.9 TFLOPS (55.6% of 34 TFLOPS). PTX tcgen05.mma does not support FP64 (documented precisions are tf32, f16/bf16, i8/u8, and f4/f6/f8), so FP64 DGEMM uses a different execution path. The additional 45% efficiency improvement is attributed to doubled FP64 units and improved memory access coalescing.
@@ VII-C2 Sustained Memory Bandwidth
Memory-intensive scientific applications require sustained high-bandwidth data movement capabilities. We employ the STREAM Triad benchmark to measure achievable memory bandwidth. Table XIII reports results from our B200 runs. Triad uses three arrays, so 64 GB and 128 GB array sizes would require device memory 192 GB and 384 GB respectively; our evaluation node had insufficient memory for those sizes, so we report only 4–16 GB.
| Array Size | B200 (TB/s) | B200 % peak |
| 4 GB | 4.141 | 51.8 |
| 16 GB | 4.140 | 51.8 |
At 4–16 GB working sets, B200 achieves 4.14 TB/s (52% of 8 TB/s peak). Larger arrays would be needed to approach peak bandwidth but require devices with 192 GB memory for 64 GB and 384 GB for 128 GB in our benchmark setup.
@@ VII-C3 Sparse Operations
| webbase-1M | 99.99% | 5.08 | 3.16× | 39.32 |
| circuit5M | 99.95% | 4.94 | 3.16× | 201.44 |
| ldoor | 99.98% | 5.03 | 3.16× | 71.93 |
Using compressed sparse representations (software compression), SpMV achieves an estimated 3.16 speedup over an uncompressed baseline in our benchmarks; run-length encoding (RLE) yields approximately 8.2 compression ratio for sparse row pointer arrays. Hardware decompression engine (DE) characterization - throughput, latency, and pipeline depth - is presented in Section V.
@@ VII-D Mixed-Precision Training: End-to-End Training Performance
We present comprehensive training benchmarks across different model architectures to assess the practical impact of architectural improvements in realistic training scenarios, as summarized in Table XI. Training speedup decomposes into SM count (1.09), CTA pairing (1.27), and TMEM (1.26). Energy efficiency improves for GPT training despite higher power consumption.
@@ VIII Discussion
Table X provides a comprehensive overview of performance improvements across all evaluated workload categories, highlighting the specific architectural features responsible for each performance gain.
Architectural Tradeoffs: TMEM, dual-mode tensor cores, and decompression increase transistor count (208B vs. 180B) but deliver 1.5–3.9 gains. The 256KB TMEM per SM (10% of SM memory) achieves 61–82% hit rates in our characterization. Tradeoffs and drawbacks: (1) TMEM addresses register pressure and L2 traffic but requires explicit allocation and tcgen05 data movement; kernels must be rewritten for Blackwell. (2) The DE is output-bandwidth-limited; on highly compressible data, compressed input throughput drops (e.g., scaling), so DE use is most beneficial when decompressed output, not input bandwidth, is the bottleneck. (3) PTX tcgen05.mma does not support FP64; FP64 DGEMM uses a different path (e.g., doubled FP64 units), so TMEM does not directly improve FP64 scientific kernels.
Software Ecosystem: CUDA 13.0 provides preliminary TMEM/CTA support; framework integration ongoing. FP6 hardware support exists but lacks software tooling. FP4/FP6 require per-layer precision selection—8.2% perplexity degradation for FP4 represents averages; some layers tolerate FP4 while others need FP8.
Performance guidelines. We consolidate actionable recommendations: (a) Use TMEM for accumulator staging and multi-stage tensor pipelines (e.g., fused attention then softmax then value) when working sets exceed L1; for single-shot small matrices, minimize TMEM round-trips. (b) Scale DE usage with compression ratio: expect 170–220 GB/s decompressed output; compressed input rate scales inversely with ratio. (c) Leverage warp-level tcgen05.mma and CTA-pair execution for GEMM and LLM inference; target 6464 tiles for TMEM. (d) For LLM inference, FP8 balances throughput and quality; FP4 yields higher throughput with acceptable perplexity increase when validated per model.
Deployment: For LLM inference, B200 provides 1.8–3.9 advantages; FP4 practical for 70B models. Training improvements (1.85 ResNet-50, 1.55 GPT-1.3B) enable larger batches. HPC gains (1.92 FP64) competitive for scientific computing.
@@ IX Conclusion
NVIDIA’s B200 GPU marks a major shift in GPU architectures. Our work presents the first detailed microbenchmark suite-based characterization of the NVIDIA Blackwell B200 GPU. Our work offers insights into its architectural innovations and performance behavior. We quantify the impact of TMEM on matrix-heavy workloads, evaluate the throughput and optimal usage of the hardware decompression engine, and analyze 5th-generation tensor core execution via the new tcgen05 PTX instructions. Our study further assesses FP4 and FP6 precision trade-offs, benchmarks Blackwell across diverse workloads—including LLM inference, scientific kernels, and mixed-precision training—and distills actionable performance guidelines for developers targeting this next-generation architecture. Our experiments primarily exercised single-die behavior; dual-die (NV-HBI) interactions and cross-die latency remain as future work.
Acknowledgment
This research used resources from the NVIDIA Brev Cloud Compute and the Google Cloud Provider. This material is based upon work supported by the U.S. DOE under Contract DE-FOA-0003177, S4PST: Next Generation Science Software Technologies Project.